# No-prior Bayesian inference reIMagined: probabilistic approximations of inferential models

### v1

@@ 2.2 Possibilistic IMs
The relative likelihood defines a possibility contour, i.e., a non-negative function such that $\sup_{\theta}R(x,\theta)=1$ for almost all $x$. This contour determines a possibility measure that can be used for data-driven uncertainty quantification about $\Theta$, which has been extensively studied in (e.g., Shafer, 1982; Wasserman, 1990a ; Denœux, 2006, 2014). This likelihood-driven possibility has a number of desirable properties. What it lacks, however, is a justification for why the “possibilities” assigned to hypotheses about $\Theta$ have belief-forming inferential weight. Fortunately, it’s straightforward to achieve the desired goal by “validifying” (Martin, 2022a ) the relative likelihood; see, also, Martin, (2015, 2018). This amounts to applying a version of the probability-to-possibility transform (e.g., Dubois et al., 2004; Hose, 2022), and the result is the possibilistic IM’s contour function:
@@
An illustration of what a typical contour $\pi_{x}$ looks like is shown in Figure 1.
@@ 3 Inner probabilistic approximations
Such a $\mathsf{Q}_{x}^{\star}$ is called the inner probabilistic approximation of the possibility measure $\overline{\Pi}_{x}$ (Dubois et al., 2004). Note that $C_{\alpha}(x)$ in (5) is a $100(1-\alpha)$credible set associated with an inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ and, since the IM’s validity property implies that it’s also a $100(1-\alpha)$confidence set, it follows that $\mathsf{Q}_{x}^{\star}$ exactly—not just asymptotically—achieves the “probability matching” property (e.g., Staicu and Reid, 2008; Datta and Ghosh, 1995) desired of no-prior Bayes solutions. That is, the $\mathsf{Q}_{x}^{\star}$-credible sets are also confidence sets, by definition/construction.
What does it take to satisfy (9)? From the characterization (8) in Theorem 1, it is relatively easy to see that equality in (9) amounts to a choice of kernel $\mathsf{K}_{x}^{\beta}$ that is fully supported on $\partial C_{\beta}(x)$, the boundary of $C_{\beta}(x)$, and marginal $\mathsf{M}_{x}={\sf Unif}(0,1)$. How the kernel allocates its mass to $\partial C_{\beta}(x)$ isn’t determined by the equality in (9), hence the inner probabilistic approximation of $\overline{\Pi}_{x}$ isn’t unique. But the sets $C_{\beta}(x)$ are typically bounded, i.e., contained in a compact subset of $\mathbb{T}$; in fact, when the sample size is moderate to large, they’re approximately ellipsoids (Section 4.2). So, in such cases, choosing the kernel $\mathsf{K}_{x}^{\beta}$ to be a uniform distribution on $\partial C_{\beta}(x)$ is quite natural—and consistent with both the maximum entropy principle (e.g., Cover and Thomas, 2006; Bernardo and Smith, 1994) and the indifference principle (e.g., Keynes, 1921; Jaynes, 2003).
@@
    sample $(\Theta\mid x,A)\sim\mathsf{K}_{x}^{A}:={\sf Unif}\{\partial C_{A}(x)\}$.
Step 2 is far simpler to state than it is to actually carry out, and I’ll discuss how this challenge can be addressed—at least approximately—in Section 5.
@@ 4.1 Agrees with Bayes in group invariant models
For example, if the distribution of $X$ depends on a location parameter $\theta$, then the distribution of $X+\tau$ depends on parameter $\theta+\tau$. When the statistical model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ satisfies (10), it’s called an invariant statistical model. Let $\overline{\mathscr{G}}$ denote the collection of transformations $\bar{g}$ on $\mathbb{T}$, corresponding to the mappings $g\in\mathscr{G}$ and parameters $\theta\in\mathbb{T}$; it’s easy to check that $\overline{\mathscr{G}}$ is a group too. Much more details on group invariant models can be found in Fraser, (1968), Eaton, (1989), and Schervish, (1995, Ch. 6), among others. The necessary technical background is given in Appendix B.
When the model is invariant in the sense described above, there are standard and broadly agreed upon no-prior Bayes solutions, namely, those based on combining the likelihood with the so-called right Haar prior using Bayes’s formula (e.g., Kass and Wasserman, 1996). This also agrees with the solution obtained via Fisher’s fiducial argument, among others. In Martin, 2023a , I showed that, under the above conditions, there is a connection between the possibilistic IM and this standard no-prior Bayes solution. This was before the inner probabilistic approximation was available more generally, so my goal there was different than it is here. But the result established there states that the standard no-prior Bayes solution—namely, the Bayes posterior with respect to the invariant right Haar prior—is precisely the inner probabilistic approximation of the possibilistic IM. Further details, including a new proof, are provided in Appendix B.
@@
For an invariant statistical model, under the conditions laid out in Appendix B, the Bayesian posterior distribution relative to the right Haar prior is the inner probabilistic approximation of the possibilistic IM.
That is, under an invariant statistical model, the familiar Bayes/fiducial distribution is the inner probabilistic approximation of the possibilistic IM. This connection serves as a strong proof of concept for the new and general no-prior Bayes solution that I’m proposing here. It also helps to explain certain phenomena from the Bayesian perspective. For example, this inner probabilistic approximation connection explains why the Bayesian/fiducial solutions in these problems are exactly probability matching.
@@ 4.2 Asymptotic normality and efficiency
where $\mathcal{U}$ is a fixed but arbitrary compact subset of $\mathbb{R}$, and $\gamma$ is the standard Gaussian possibility with zero mean and identity covariance matrix.
@@
In light of Theorem 3, the IM’s inner probabilistic approximation converges to a Gaussian probability distribution just like in the Bayesian Bernstein–von Mises theorem. Here I’ll give a quick sketch of the proof for a scalar parameter $\Theta$. For a generic value $\theta$, if the level set $C_{\alpha}(x^{n})$ is the interval $[\text{low}_{\alpha}(x^{n}),\text{upp}_{\alpha}(x^{n})]$, then it’s easy to verify that
$\{\text{low}_{\alpha}(x^{n})\geq\theta\}\;\text{or}\;\{\text{upp}_{\alpha}(x^{n})\leq\theta\}\iff\pi_{x^{n}}(\theta)\leq\alpha.$
@@
Similarly, if $z>0$, then $1-\tfrac{1}{2}\,\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}z)\to\mathsf{P}(Z\leq z)$. Therefore, with probability tending to 1, the distribution of $J_{X^{n}}^{1/2}(\Theta-\hat{\theta}_{X^{n}})$, under $\Theta\sim\mathsf{Q}_{X^{n}}^{\star}$, is converging to a standard normal, just like the classical Bernstein–von Mises theorem says.
@@ 4.3 Marginalization risks, or lack thereof
That the possibility-theoretic marginalization strategy appears organically in the common-sense treatment of confidence sets is no coincidence; see Martin, 2021a . The inner probabilistic approximations of $\overline{\Pi}_{x}$ are confidence distributions, and these are characterized as mixtures of kernels supported on the boundaries of the confidence sets. Then the derived marginal distribution $\mathsf{Q}_{x}^{m}$ of $\Phi=m(\Theta)$ clearly also can be characterized by a similar mixture, but $\mathsf{Q}_{x}^{m}$ would only be the inner probabilistic approximation of $\overline{\Pi}_{x}^{m}$ if the boundaries $\partial C_{\alpha}(x)$ of the original confidence sets get mapped via $m$ to the boundaries $\partial C_{\alpha}^{m}(x)$ of the marginal confidence sets. Define $m$ to be boundary-preserving if $m\{\partial C_{\alpha}(x)\}=\partial C_{\alpha}^{m}(x)$ for each $\alpha\in[0,1]$. Among other forms of mappings, linear transformations are boundary preserving because they attain their extrema on the boundary of their domain.
@@
That the probability calculus leads to reliable marginalization effectively only when the transformation is linear doesn’t imply that reliable marginalization is out of reach more generally. It just means non-probabilistic marginalization strategies are needed. This isn’t unfamiliar: Bayesians choose default prior distributions that are tailored to the particular inferential target $\Phi$ (e.g., Tibshirani, 1989; Datta and Ghosh, 1995; Bernardo, 1979; Berger et al., 2009), and confidence and fiducial distribution users carry out various ad hoc manipulations akin to those used to construct approximate pivotal quantities. My new suggestion is to first carry out a more efficient brand of marginalization on the possibilistic IM, one based on profiling, and then obtain the inner probabilistic approximation of that marginal possibilistic IM. Details of this will be presented in the context of a concrete example in Section 6 below.

### v2

# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

@@ 2.2 Possibilistic IMs
The relative likelihood defines a possibility contour, i.e., a non-negative function such that $\sup_{\theta}R(x,\theta)=1$ for almost all $x$. This contour determines a possibility measure that can be used for data-driven uncertainty quantification about $\Theta$, which has been extensively studied in (e.g., Shafer, 1982; Wasserman, 1990a ; Denœux, 2006, 2014). This likelihood-driven possibility has a number of desirable properties. What it lacks, however, is a justification for why the “possibilities” assigned to hypotheses about $\Theta$ have belief-forming inferential weight. Fortunately, it’s straightforward to achieve the desired goal by “validifying” (Martin, 2022a ; Martin, 2022b ) the relative likelihood; see, also, Martin, (2015, 2018). This amounts to applying a version of the probability-to-possibility transform (e.g., Dubois et al., 2004; Hose, 2022), and the result is the possibilistic IM’s contour function:
@@
An illustration of what a typical contour $\pi_{x}$ looks like is shown in Figure 1. Some further details about this simple example are presented at the end of Section 3 below.
@@ 3 Inner probabilistic approximations
There are cases in which attainment of the equality in (9) is out of reach. For instance, suppose that $\theta\mapsto\pi_{x}(\theta)$ is bounded away from 0 for some $x$. This is rare in applications, but not impossible. It may happen that the data is not particularly informative about a feature $\Phi=m(\Theta)$ of $\Theta$, and then the corresponding marginal IM (see Section 4.3) has a non-vanishing contour, i.e., the data is insufficiently informative to exclude any values of $\Phi$ with high degree of confidence. In any case, a probability distribution $\mathsf{Q}_{x}^{\star}$ satisfying (9), if one exists, will be called the inner probabilistic approximation of the possibility measure $\overline{\Pi}_{x}$ (Dubois et al., 2004).
Note that $C_{\alpha}(x)$ in (5) is a $100(1-\alpha)$credible set associated with an inner probabilistic approximation $\mathsf{Q}_{x}^{\star}$ and, since the IM’s validity property implies that it’s also a $100(1-\alpha)$confidence set, it follows that $\mathsf{Q}_{x}^{\star}$ exactly—not just asymptotically—achieves the “probability matching” property (e.g., Staicu and Reid, 2008; Datta and Ghosh, 1995) desired of no-prior Bayes solutions.
Theorem 2.
$\mathsf{Q}_{x}^{\star}$-credible sets are confidence sets.
What does it take to satisfy (9)? From the characterization (8) in Theorem 1, it is relatively easy to see that equality in (9) amounts to a choice of kernel $\mathsf{K}_{x}^{\beta}$ that is fully supported on $\partial C_{\beta}(x)$, the boundary of $C_{\beta}(x)$, and marginal $\mathsf{M}_{x}={\sf Unif}(0,1)$. How the kernel allocates its mass to $\partial C_{\beta}(x)$ isn’t determined by the equality in (9), hence the inner probabilistic approximation of $\overline{\Pi}_{x}$ isn’t unique. But the sets $C_{\beta}(x)$ are typically bounded, i.e., contained in a compact subset of $\mathbb{T}$, and in fact, when the sample size is moderate to large, they’re approximately ellipsoids (Section 4.2). So, in such cases, choosing the kernel $\mathsf{K}_{x}^{\beta}$ to be a uniform distribution on $\partial C_{\beta}(x)$, for each $\beta\in(0,1)$, is quite natural, and consistent with both the maximum entropy (e.g., Cover and Thomas, 2006; Bernardo and Smith, 1994) and indifference principles (e.g., Keynes, 1921; Jaynes, 2003).
@@
    sample $(\Theta\mid x,A=\alpha)\sim\mathsf{K}_{x}^{\alpha}:={\sf Unif}\{\partial C_{\alpha}(x)\}$.
Step 2 is far simpler to state than it is to carry out, and I’ll discuss how this computational challenge can be addressed—at least approximately—in Section 5.
As a quick illustration, consider a gamma model $\mathsf{P}_{\theta}={\sf Gamma}(n,\theta)$ for a single observation $X$, with $n$ a fixed constant; the more practical case of $n$ iid exponential samples where $X$ is the sum, the complete sufficient statistic, reduces to this one. Then the likelihood-based possibilistic IM has contour function given by
$\pi_{x}(\theta)=\mathsf{P}_{\theta}\bigl{\{}(X/\theta)^{n}e^{-X/\theta}\leq(x/\theta)^{n}e^{-x/\theta}\bigr{\}},\quad\theta>0.$
The right-hand side has no simple closed-form expression but, since $X/\theta$ is a pivot under $\mathsf{P}_{\theta}$, the computations are straightforward using Monte Carlo; a complicated exact expression, whose evaluation doesn’t require Monte Carlo, is given in Proposition 1 of Hector et al., (2025). The contour shown in Figure 1 is actually the one above, with $n=7$ and $x=14$. Clearly, the level sets $C_{\alpha}(x)$ are intervals, say, $[a(x,\alpha),b(x,\alpha)]$, where $b\geq a$, but with $a(x,1)=b(x,1)=x/n$, the maximum likelihood estimator. Then the kernel $\mathsf{K}_{x}^{\alpha}$ in the inner probabilistic approximation is a probability distribution supported on the endpoints $a(x,\alpha)$ and $b(x,\alpha)$, i.e.,
$\mathsf{K}_{x}^{\alpha}=w_{x,\alpha}\,\delta_{a(x,\alpha)}+(1-w_{x,\alpha})\,\delta_{b(x,\alpha)},$
where $w_{x,\alpha}\in[0,1]$ and $\delta_{z}$ denotes a probability measure that puts all its mass at $z$. The question I want to address here is: what kind of distribution is an inner probabilistic approximation, what does it look like? In the case where $w_{x,\alpha}\equiv w$ is a fixed value that doesn’t depend on $\alpha$, the distribution function of $\Theta\sim\mathsf{Q}_{x}^{\star}$ can easily be derived:
$\mathsf{Q}_{x}^{\star}(\Theta\leq t)=\begin{cases}w\,\pi_{x}(t)&\text{if $t\leq x/n$}\\ 1-(1-w)\,\pi_{x}(t)&\text{if $t>x/n$}.\end{cases}$
Two such distribution functions are shown in Figure 2(a), one based on $w=0.5$ (solid) and another based on $w\approx 0.45$ (dashed). A third distribution function (dashed) is also shown, namely, the Bayesian posterior distribution function based on a default (right Haar) prior with density $\theta\mapsto\theta^{-1}$, which is an inverse gamma distribution. This Bayes posterior doesn’t obviously have a form like that above, [^2] I wasn’t able to find a constant $w$ that would recover the inverse gamma distribution—the choice $w\approx 0.45$ was my best effort. Consequently, this particular instantiation of the general form of $\mathsf{Q}_{x}^{\star}$ must involve an $\alpha$-dependent weight and it’d be interesting to determine what that $w_{x,\alpha}$ is. but that it too is an inner probabilistic approximation of the IM is a general fact established in Section 4.1 below. The three distribution functions are similar, but not exactly the same. To confirm that all three are, indeed, inner probabilistic approximations, I plot, in Figure 2(b), the non-coverage probability functions $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$. According to (9), the non-coverage probability for each inner probabilistic approximation should be equal to $\alpha$, which is confirmed in the plot.
(a) Distribution function of $\mathsf{Q}_{x}^{\star}$
(b) Non-coverage probability of $\mathsf{Q}_{x}^{\star}$
Figure 2: Plots related to three inner probabilistic approximations for the gamma illustration with $n=7$ and $x=14$: uniform kernel (solid), non-uniform kernel with $w\approx 0.45$ (dashed), and the Bayesian posterior distribution (dotted) based on the default prior $\theta\mapsto\theta^{-1}$. Panel (a) shows the cumulative distribution function $t\mapsto\mathsf{Q}_{x}^{\star}(\Theta\leq t)$ and Panel (b) shows the non-coverage probability function $\alpha\mapsto 1-\mathsf{Q}_{x}^{\star}\{C_{\alpha}(x)\}$.
@@ 4.1 Agrees with Bayes in group invariant models
For example, if the distribution of $X$ depends on a location parameter $\theta$, then the distribution of $X+\tau$ depends on parameter $\theta+\tau$. When the statistical model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ satisfies (10), it’s called an invariant statistical model. Let $\overline{\mathscr{G}}$ denote the collection of transformations $\bar{g}$ on $\mathbb{T}$, corresponding to the mappings $g\in\mathscr{G}$ and parameters $\theta\in\mathbb{T}$; it’s easy to check that $\overline{\mathscr{G}}$ is a group too. More details on group invariant models can be found in Fraser, (1968), Eaton, (1989), and Schervish, (1995, Ch. 6), among others. The necessary technical background is given in Appendix B.
When the model is invariant in the sense described above, there are standard and broadly agreed upon no-prior Bayes solutions, namely, those based on combining the likelihood with the so-called right Haar prior using Bayes’s formula (e.g., Kass and Wasserman, 1996). This also agrees with the solution obtained via Fisher’s fiducial argument, among others. In Martin, 2023a , I showed that, under the above conditions, there is a close connection between the possibilistic IM and this standard no-prior Bayes solution. In the present context, the natural interpetation of the result established there is that the standard no-prior Bayes solution—namely, the Bayes posterior with respect to the invariant right Haar prior—is an inner probabilistic approximation of the possibilistic IM. Further details, including a new proof, are provided in Appendix B.
@@
For an invariant statistical model, under the conditions laid out in Appendix B, the Bayesian posterior distribution relative to the right Haar prior is an inner probabilistic approximation of the possibilistic IM.
That is, under an invariant statistical model, the familiar Bayes/fiducial distribution is an inner probabilistic approximation of the possibilistic IM. This justifies my previous claim that the Bayes posterior in the gamma illustration, as shown in Figure 2 above, is an inner probabilistic approximation. This connection also serves as a strong proof of concept for the no-prior Bayes solution that I’m proposing here: that is, the right Haar prior Bayesian solutions in such problems are reasonable and widely used, so what I’m proposing must be equally reasonable and usable. Finally, it also helps to explain certain phenomena from the Bayesian perspective, namely, why the Bayesian/fiducial solutions in these problems are exactly probability matching.
A point made above that deserves to be mentioned again is that the inner probabilistic approximations are not unique. So, the Bayes solution here might not exactly correspond to the generically recommended strategy that defines $\mathsf{K}_{x}^{\beta}$ to be uniform on $\partial C_{\beta}(x)$; the illustration in Figure 2 is such a case. In fact, when a direct connection between a Bayesian solution and the IM’s inner probabilistic approximation can be made, then I might recommend the former since its computation is more familiar than the latter’s. But, like in Figure 2, the difference between these inner probabilistic approximations will be negligible and will vanish as the information in data increases; see Section 4.2.
As an illustration, consider a case involving directional data, where the observations correspond to directions—points on a unit circle in the plane—or simply angles with respect to a reference direction. Real-world applications involving these kind of data include wind and animal movement studies; see Mardia and Jupp, (2000) for a comprehensive treatment. More generally, direction measurements can be represented by points on the surface of a hypersphere. This kind of data is common in astronomy, where the position of planets, stars, etc. can be described by a point on the celestial sphere. The present example involves $n=25$ such planar directional observations, plotted in Figure 3. These data are simulated from a model very similar ot that used to analyze a very small data set on roulette wheel spins from Example 1.1 in Mardia and Jupp, (2000). The posited model is the so-called von Mises distribution, parametrized by a mean direction/angle $\theta$ and a concentration parameter $\kappa$. Further details about this model and the corresponding IM/Bayes/fiducial solutions are given in Appendix B.2. Here I’ll take the concentration parameter $\kappa=4$ to be fixed and known. I’ll also interpret the data as angles (radians) so that the parameter $\theta$ acts like a location paramter. With this simplification, the $\theta$-only von Mises model is invariant with respect to addition modulo $2\pi$. Therefore, Theorem 3 implies that the Bayes/fiducial distribution is an inner probabilistic approximation of the possibilistic IM. Figure 4 shows, first, a plot of the possibilistic IM’s contour for these data, with its peak at the maximum likelihood estimator, $\hat{\theta}_{x}=1.35$ and, second, a histogram of samples drawn from the inner probabilistic approximation (with $\mathsf{K}_{x}^{\beta}$ a uniform distribution) and the no-prior Bayesian posterior density overlaid. The similarities provided a visual confirmation of the claim in Theorem 3.
Figure 3: Plot of the directional data analyzed in Section 4.1. Points on the circle determine the direction. The small closed circle denotes the sample mean in Cartesian coordinates and the corresponding angle determined by that point, $\hat{\theta}_{x}=1.35$, is the maximum likelihood estimator of the mean angle. The length of that vector, 0.87, is a to-be-conditioned-on ancillary statistic; see Appendix B.2.
(a) Possibilistic IM contour
(b) Inner probabilistic approximation
Figure 4: Analysis of the directional data in Section 4.1. Panel (a) shows the possibilistic IM contour and Panel (b) shows a histogram of samples from the corresponding inner probabilistic approximation with the flat-prior Bayes posterior density overlaid.
@@ 4.2 Asymptotic normality and efficiency
where $\mathcal{U}$ is a fixed but arbitrary compact subset of $\mathbb{R}^{D}$, and $\gamma$ is the standard Gaussian possibility with zero mean and identity covariance matrix.
@@
A direct consequence of Theorem 4 is that the IM’s inner probabilistic approximation converges to a Gaussian probability distribution just like in the classical, Bayesian Bernstein–von Mises theorem. The proof is in Appendix C.
Theorem 5.
Under the conditions of Theorem 4, the approximation $\mathsf{Q}_{X^{n}}^{\star}$ of the possibilistic IM satisfies a probabilistic Bernstein–von Mises theorem, i.e.,
$d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}}):=\sup_{H}\bigl{|}\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)\bigr{|}\to 0\quad\text{in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$},$
where $\mathsf{G}_{X^{n}}={\sf N}_{D}(\Theta+n^{-1/2}\Delta_{\Theta}(X^{n}),(nI_{\Theta})^{-1})$ is the approximating Gaussian distribution and the supremum is over all Borel measurable subsets $H$ of $\mathbb{T}$.
I’ll give a quick sketch of the proof for a scalar parameter $\Theta$. For a generic value $\theta$, if the level set $C_{\alpha}(x^{n})$ is the interval $[a_{\alpha}(x^{n}),b_{\alpha}(x^{n})]$, then it’s easy to verify that
$\{a_{\alpha}(x^{n})\geq\theta\}\;\text{or}\;\{b_{\alpha}(x^{n})\leq\theta\}\iff\pi_{x^{n}}(\theta)\leq\alpha.$
@@
Similarly, if $z>0$, then $1-\tfrac{1}{2}\,\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}z)\to\mathsf{P}(Z\leq z)$. Therefore, with probability tending to 1, the distribution of $J_{X^{n}}^{1/2}(\Theta-\hat{\theta}_{X^{n}})$, under $\Theta\sim\mathsf{Q}_{X^{n}}^{\star}$, is converging to a standard normal, identical to the classical Bernstein–von Mises theorem.
A point that follows immediately from the above sketch is that assigning probability $\frac{1}{2}$ to the boundary points in the inner probabilistic approximation is crucial; of course, any choice that assigns probability to the boundary points that converges to $\frac{1}{2}$ works fine too. More generally, since $\mathsf{G}_{X^{n}}$ is an inner probabilistic approximation to the possibility measure with contour $\gamma_{X^{n}}$, and corresponds to taking a uniform kernel on the level sets of $\gamma_{X^{n}}$, it’s necessary that the choice of kernel defining $\mathsf{Q}_{X^{n}}^{\star}$ at least be asymptotically uniform. This implies that the inner probabilistic approximation’s non-uniqueness vanishes as $n\to\infty$. This observation adds further support for my recommendation above to adopt the uniform kernel in the inner probabilistic approximation.
@@ 4.3 Marginalization risks, or lack thereof
That the possibilistic marginalization strategy appears organically in the common-sense treatment of confidence sets is no coincidence; see Martin, 2021a . The inner probabilistic approximations of $\overline{\Pi}_{x}$ are confidence distributions, and these are characterized as mixtures of kernels supported on the boundaries of the confidence sets. Then the derived marginal distribution $\mathsf{Q}_{x}^{m}$ of $\Phi=m(\Theta)$ clearly also can be characterized by a similar mixture, but $\mathsf{Q}_{x}^{m}$ would only be the inner probabilistic approximation of $\overline{\Pi}_{x}^{m}$ if the boundaries $\partial C_{\alpha}(x)$ of the original confidence sets get mapped via $m$ to the boundaries $\partial C_{\alpha}^{m}(x)$ of the marginal confidence sets. Define $m$ to be boundary-preserving if $m\{\partial C_{\alpha}(x)\}=\partial C_{\alpha}^{m}(x)$ for each $\alpha\in[0,1]$. Among other forms of mappings, linear transformations are boundary preserving because they attain their extrema on the boundary of their domain.
@@
That the probability calculus leads to reliable marginalization effectively only when the transformation is linear doesn’t imply that reliable marginalization is out of reach in the non-linear case. It just means non-probabilistic marginalization strategies are needed. This isn’t unfamiliar: Bayesians choose default prior distributions that are tailored to the particular inferential target $\Phi$ (e.g., Tibshirani, 1989; Datta and Ghosh, 1995; Bernardo, 1979; Berger et al., 2009), and confidence and fiducial distribution users carry out various manipulations akin to those used to construct approximate pivotal quantities. The above insights motivate the following suggestion: first carry out the appropriate possibilistic marginalization to get $\pi_{x}^{m}$ and $\overline{\Pi}_{x}^{m}$, and then construct the inner probabilistic approximation of this. An example of this using the extension-based marginalization is shown below. Extension-based marginalization is a simple and very general, but is not the most efficient. Typically, sharper inferences can be drawn when marginalization is carried out by “validifying” a relative profile likelihood. Details of this profile-based approach will be presented in the context of a concrete example in Section 6 below.
Here I’ll revisit the directional data example presented in Section 4.1 above, where the quantity of interest is $\Phi=\cos(a\Theta)$, for two different values of $a$. Start with the case of $a=1$. Since the Bayes/fiducial distribution—which is also an inner probabilistic approximation of the possibilistic IM by Theorem 3—is almost fully supported on the interval $(0,\pi)$, where $\theta\mapsto\cos(\theta)$ is one-to-one, this mapping is effectively boundary-preserving. Then Theorem 6 implies that the marginal Bayes/fiducial distribution for $\Phi$ is an inner probabilistic approximation of the marginal possibilistic IM for $\Phi$. Figure 5 shows, for $a=1$, the extension-based marginal IM contour for $\Phi$, samples of the marginal Bayes/fiducial with the corresponding density function overlaid, and the non-credibility function $\alpha\mapsto 1-\mathsf{Q}_{x}^{m}\{C_{\alpha}^{m}(x)\}$, for $\mathsf{Q}_{x}^{m}$ equal to the Bayes/fiducial marginal distribution and to the inner probabilistic approximation of the extension-based marginal IM. The key observation is that the contour and the posterior distribution have similar shapes and, as Theorem 6 predicts, the non-credibility function behaves exactly as it should for an inner probabilistic approximation, according to the definition (9). Again, this all happens because $\theta\mapsto\cos(a\theta)$ with $a=1$ is approximately boundary-preserving on the interval where the relevant posterior distributions are supported.
(a) Marginal IM contour
(b) Bayes marginal
(c) Non-credibility
Figure 5: Marginal inference results for $\Phi=\cos(a\Theta)$ with $a=1$ in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).
Next, take $a=1.5$. Then the aforementioned “effectively boundary-preserving” property no longer holds and, surprisingly, this seemingly small change in the quantity of interest has non-negligible impact on the results. Figure 6 shows the same plots as in Figure 5, this time for $a=1.5$. The key observation is that, while the general shape of the marginal IM contour and the Bayes posterior distribution were similar in Figure 5, they are quite different in Figure 6; this is a consequence of the entirely different style of marginalization, from $\Theta$ to $\Phi$, with the IM’s possibilistic marginalization preserving the validity properties enjoyed by both approaches’ inference on $\Theta$ whereas the Bayesians’ probabilistic marginalization doesn’t. The IM’s inner probabilistic approximation (not shown) looks similar to the Bayesian posterior, but it’s not the same. Indeed, Figure 6(c) shows the non-credibility for the Bayes posterior distribution and the IM’s inner probabilistic approximation, and there are clear signs of the Bayesians’ over-confidence. The IM’s inner approximation also shows (milder) signs of over-confidence, but there’s a perfectly reasonable explanation for this. Recall from Section 3 that exact inner probabilistic approximations need not exist, e.g., if the possibility contour is bounded away from 0. In Figure 6(a), note that the contour is in fact bounded away from 0 and, consequently, the sets $C_{\alpha}^{m}(x)$ are equal to the full space $[-1,1]$ for sufficiently small $\alpha$. Since any probability distribution $\mathsf{Q}_{x}^{m}$ supported on $[-1,1]$ must assign probability $1>1-\alpha$ to $C_{\alpha}^{m}(x)=[-1,1]$ for all sufficiently small $\alpha$, there is no $\mathsf{Q}_{x}^{\star}$ that satisfies (9) for all $\alpha$. But the IM’s pseudo inner probabilistic approximation—“pseudo” because the inner probabilistic approximation strictly doesn’t exist—quickly corrects for this over-confidence at small $\alpha$ and achieves the objective (9) for all $\alpha$ that are not too small.
(a) Marginal IM contour
(b) Bayes marginal
(c) Non-credibility
Figure 6: Marginal inference results for $\Phi=\cos(a\Theta)$ with $a=1.5$ in the directional data example. Panel (a) shows the extension-based marginal IM contour for $\Phi$; Panel (b) shows the Bayesian marginal posterior distribution for $\Phi$; and Panel (c) shows the non-credibility function for the marginal Bayesian posterior (dashed) and the marginal IM’s inner probabilistic approximation (solid).